# The Riemann Hypothesis is false

### v42

@@
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}$ (9)
say. From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
Let $J(u)=\sum_{0<r\leq u}e^{2\pi if(r)}=\sum_{0<r\leq M}e^{2\pi if(r)}$ where $u\leq m\varepsilon,M=M(u)$ is the largest integer $\leq u$, $f(r)=\frac{T\log(1+\frac{r}{m})}{2\pi}$ and $m\in\mathbb{N}$. Then $f^{\prime}(r)=\frac{T}{2\pi(r+m)}$. Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and choose $T\in\mathbb{Q}$ so that both $\alpha$ and $\beta$ are not integers. Note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$. Thus if $\eta$ is any constant $\in(0,\frac{1}{2})$, then the range $[\alpha-\eta,\beta+\eta]$ doesn’t contain any integer. Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0$ and by Lemma 4.7 of Titchmarsh [7] we have
@@
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log m).$ (23)
@@
$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log T).$ (29)
@@
$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log T).$ (32)
@@
$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}(\log T-\log(T/2\pi))+\frac{T}{2}+o(T)$ (36)
@@
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}-\frac{T}{\pi}\log(2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
since $H=2T+O(1)$. Combining (42) and (39) gives
@@
$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T)$ (46)
$=-\frac{T}{\pi}\log\log T+(1-B)\frac{T}{2\pi}+\frac{T}{2\pi}\log(2\pi)+o(T).$ (47)
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (48)
uniformly for $2\leq n\ll T$. Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3]. Thus, summing (48) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (49)
$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T).$ (50)
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n<T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (51)
$=\frac{T}{2\pi}\log(4\pi)+o(T).$ (52)

### v43

@@
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
@@
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
@@
Let $J(u)=\sum_{0<r\leq u}e^{2\pi if(r)}=\sum_{0<r\leq M}e^{2\pi if(r)}$ where $u\leq m\varepsilon,M=M(u)$ is the largest integer $\leq u$, $f(r)=\frac{T\log(1+\frac{r}{m})}{2\pi}$ and $m\in\mathbb{N}$. Then $f^{\prime}(r)=\frac{T}{2\pi(r+m)}$. Put $\alpha=f^{\prime}(M),\beta=f^{\prime}(0)$ and note that $\beta-\alpha\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$ uniformly for $0\leq u\leq m\varepsilon$. Thus if $\eta$ is any constant $\in(0,1)$, then the range $[\alpha-\eta,\beta+\eta]$ contains at most two integers. By Lemma 4.2 of Titchmarsh [7], we have $\Big{|}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\Big{|}\leq\frac{4}{f^{\prime}(M)}\ll 1.$ Thus $\sum_{\alpha-\eta\leq v\leq\beta+\eta}\int_{0}^{u}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll 1$ and by Lemma 4.7 of the same book we have
@@
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
@@
$=\frac{1}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}+O(T^{3/4}\log^{6}T).$ (29)
@@
$=\frac{2\pi i}{e}\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O(T^{-1/4})\Big{)}+O(T^{3/4}\log^{6}T).$ (32)
@@
$=(B-1)\frac{T}{2\pi}-\frac{T}{2\pi}(\log T-\log(T/2\pi)+O(\varepsilon))+\frac{T}{2}+o(T)$ (36)
@@
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{3T}{2\pi}-\frac{3T}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
since $H=3T+O(1)$. Combining (42) and (39) gives
@@
$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
Summing (9) for $\frac{T}{\log^{2}T}<n\leq T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)$ (46)
$=-\frac{2T}{\pi}\log\log T+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T).$ (47)
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{\pi n}+O((\log T)\log\log 3n)$ (48)
uniformly for $2\leq n\ll T$. Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3]. Thus, $N(H)-N(T)=N(3T)-N(T)+O(\log T)=(T/\pi)\log(T/2\pi)+(T/2\pi)(-2+3\log 3)+O(\log T)$ and $(T/\pi)\sum_{2\leq m\leq\frac{T}{\log^{2}T}}\frac{\Lambda(m)}{m}=(T/\pi)\log T-(2T/\pi)\log\log T-(B/\pi)T+o(T).$ Hence, summing (48) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (49)
$=\frac{2T}{\pi}\log\log T+\frac{T}{2\pi}(3\log 3-2\log(2\pi)+2B-2)+o(T).$ (50)
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (51)
$\gg T.$ (52)